# Testing stimulated emission photon directions

### v12

@@ I Introduction
Figure 4: In commercially available e.g. erbium-doped optical fiber amplifiers (EDFA, or simpler semiconductor SOA, etc.) [11] there is observed strong backward ASE (amplified spontaneous emission) - increased emission from pumped doped fiber toward laser [12], exactly as in the sought CPT analog of ”laser causes excitation” effect, usually removed by forward optical isolators. Top: Analogously to STED-like test, we can verify this photon direction by measuring delay between the impulse and reduced population of excited atoms of external target, e.g. by monitoring its spontaneous emission. For this purpose, the shown standard forward optical isolator should be removed, or better rotated to backward position. Bottom: simplified setting from [12] seems much more practical for tests: of delay (opposite if CPT analog), separation (leaving forward photons in CPT perspective), transparency (blocked by excited/deexcited?) and potential future applications. As in Fig. 5, it seems crucial to use reverse bias for the laser, in practice (e.g. [12, 1, 2]) probably included in varying/impulse signal.
@@
Figure 5: Usually system has tendency to get to the ground state, emitting photon, which e.g. excites some atom, according to absorption equation. However, getting reversed tendency should increase amplitude (probability) of shown symmetric reversed Feynman diagram: causing absorption of photon, emitted by some atom - according to symmetric stimulated emission equation. Beside previous based on synchrotron radiation, diagram shows two examples of obtaining such revered tendency: upper uses membrane GFP (green fluorescent protein) which can act as light-driven protein pump [17], in reversed proton gradient would get tendency to absorb photons, what might be used e.g. by biological organisms for thermoregulation. The bottom is analogously applied reversed bias/potential/electron gradient for light/laser diode, what is also applied for light detectors[18], but should actually also stimulate emission. Impulse laser sources often have e.g. sinc-like shape, so should also include reversed bias, what might explain observations in [1, 2]. Backward ASE [12] probably also uses both forward and reverse bias in varying signal, if successful would be perfect for presented applications. Also Fig. 2, 4 tests should rather use reverse bias. Alternatively, change of $N_{2}$ above/below thermal $N_{2}\sim N_{1}\exp(-\beta(E_{2}-E_{1}))$ of Boltzmann distribution should also increase amplitude of Feynman diagram with emission/absorption there, and absortpion/emission in the target.
@@ II-B Alternative STED-like tests (needs reverse bias)
Figure 8: Top: ring laser uses optical isolator to enforce unidirectional photon circulation, which would be reversed in CPT perspective for reversed bias - switching the action of absorption and stimulated emission equations, suggesting only one of them should act on external targets - let us refer to them as forward/backward beam. Analogously for just adding optical isolator, or synchrotron sources like synchrotron, betatron, wiggler, undulator, free electron laser, bremsstrahlung - reversing electron and photon trajectory in CPT perspective. FEL reaches narrower spectrum, better coherence, narrower angle [15] - what already seems a consequence of discussed CPT analog of emission: that further electrons stimulate emission from earlier (to be verified).
Center: STED microscopy (image source) showing that, at least for diode lasers, both equations act on external target, hence we can call it bidirectional beam. We usually focus on absorption by external target, because stimulated emission requires its initial excitation, while without such preparation usually $N_{2}/N\approx 0$.
Bottom: while CPT symmetry of physics is required by local Lorentz-invariant QFT, e.g. switching forward/backward beams, in theory general relativity allows its realization by travelling through nonorientable wormhole [27], bringing tangible thought experiment valuable for intuitions, e.g. switching sign of radiation pressure.
@@ IV Some potential new applications
The above considerations could lead to many new applications, especially if generating backward beam of negative radiation pressure acting on target only with stimulated emission equation. Optimistically it can be obtained by reverse biased diode/laser (without or with reversed optical isolator). Otherwise, as in Fig. 3, 8, should be possible with antennas and synchrotron radiation, e.g. betatron, synchrotron, free electron laser, or bremsstrahlung e.g. in X-ray lamp in strong magnetic field. For example:
@@
    Figure 16: Potential applications of proposed backward camera to obtain complementary information from astronomical objects, e.g. might allow detection/observation of black holes as in Fig. 17, 20, 21, maybe also finding some novel currently unexpected objects, and mapping emission asymmetry for cosmological studies. Many of such objects e.g. pulsars (image source) have accelerating charged particles, what leads to synchrotron radiation - e.g. causing excitation in cooled detector of standard telescopes. In CPT perspective of this scenario, there is still accelerating charge (in opposite direction) - should emit photons, causing excitation now toward our negative time, which is deexcitation toward positive time. Therefore, such process required by CPT symmetry would need excited telescope sensor - usually they are cooled instead, what might be the reason for not observing such effect, could lead to worth searching consequences like cooled halos around such pulsars. To have a chance to observe such effect, we would need to pump the sensor/detector (instead of standard cooling), again monitoring its population, or maybe using semi-transparent mirror as in Fig. 13 with various photon sources (e.g. antenna, synchrotron, X-ray lamp, forward bias diode with monitored current), maybe also include some optical isolator to allow only outgoing photons. If successful, such telescope should see (reversed) synchrotron, but not thermal radiation - e.g. allowing to separate them when combining with standard picture, also for observations/diagnostics/early warnings for our Sun [52].
@@
    Figure 17: Applying T/CPT symmetry to our black hole (running movie backward), it is a white hole, and they actively emit - acting with positive radiation pressure $\vec{p}=\langle\vec{E}\times\vec{H}\rangle/c$, heating matter around, what returning to our time perspective means that black hole should also actively cool matter around with negative radiation pressure. Hence white/black holes should repel/attract not only gravitationally, but also EM by radiation pressure - continuously interacting/exchanging information with the surrounding in symmetric way, this way allowing unitary evolution, avoiding information paradox [53] symmetrically for both white and black hole. It also would help forming dynamical outward/inward jets for both, and indeed they turn out very dynamical [54]. As white holes would heat matter around, black hole should symmetrically cool it - with the same spectrum, but replacing absorption with stimulated emission equation, and indeed matter [55], especially electrons [56] around turn out colder than expected, it also might prevent observation of IMBH (intermediate mass black holes). Such pull not only in gravity, but also in EM, would additionally increase growth of black hole, which indeed turns our faster than expected [57], generally might be faster than Hawking radiation - suggesting to reconsider their future evolution. While we might already observe indirect consequences of such stimulated emission caused by black hole, backward telescope like in Fig. 13, 16, 20 might also observe it directly - also below horizon, symmetrically to white hole observation. Black holes are hypothesized to have negative pressure [58], negative temperature pressure [59], which should also lead to radiation pressure through reversed emission/absorbtion tendency as in Fig. 5.
@@
    Figure 20: Potential observation of black holes with (radio)telescope. From CPT perspective: running movie backward, this would be white hole emitting photons (from below horizon) which could be absorbed by telescope - what in our time perspective means emission by telescope stimulated by black hole, however, it requires earlier excitation of telescope sensor. While usually radio telescope amplifiers are cooled, avoiding it or actively exciting resonator inside, in theory this antenna could also emit directionally. However, used frequencies are often difficult to absorb in cosmos, without specific absorber might go to infinity, what in CPT perspective would mean such EM wave was never emitted, hence emission should be forbidden if not pointing absorber, like black hole - suggesting novel way to detect them, and observe also below horizon, symmetrically to white hole emitting from below horizon. Another view is by $\langle\psi_{f}|U|\psi_{i}\rangle$ S-matrix: with amplitude/probability of photon depending both on emitter in $\psi_{i}$, but also absorber in $\psi_{f}$, they are switched in CPT view - without absorber, in CPT view there is no emitter.
    Figure 21: Wheeler-Feynman absorber theory requires both standard retarded EM waves, but also symmetrically advanced, and they are switched in CPT view. As general relativity is solved by the least action principle in Einstein’s field equations, gravitational waves should also propagate symmetrically. E.g. LIGO measures length, which does not change under T/CPT symmetry, hence should observe both retarded and advanced waves, the question is how to distinguish them. One way is comparison with separate e.g. retarded visual observation - its lack could indicate advanced wave, which might be also suggested by delays between detectors. Diagram shows example of such hypothetical event: assuming Big Crunch and evolving backward as in Fig. 18, symmetry suggests there should be also tendency to form black holes which can merge, leading to spacetime distortion with advanced wave propagation, viewed by e.g. LIGO as chirp of time-reversed chirp. It suggests to just search historical LIGO data for standard dataset of chirps, but with applied time-symmetry to reverse their shape. Finding such event could be used as confirmation of Big Crunch hypothesis. While LIGO just measuring lengths should see both retarded and advanced waves, EM receivers are asymmetric: usually focused on absorption - of retarded waves, for advanced would need initial excitation to be sensitive for stimulated emission.
@@
  • Backward camera as in Fig. 13 maybe exploiting transparency and/or opposite delay e.g. for astronomical observations (of e.g. synchrotron radiation of pulsars, our Sun [52], maybe black holes as in Fig. 16, 17)- focused on target’s negative radiation pressure (or transmission of added backward beam), while standard (forward) camera is focused on positive. It would replace CMOS/CCD matrix behind the lens, with matrix/surface of continuously excited e.g. approximately 3-level medium like dye or erbium, plus some monitoring of its population map - e.g. through spontaneous emission, or EIT (electromagnetically induced transparency), half-silvered mirror and CMOS/CCD, or maybe forward biased diode with monitored current. Pairing with backward beam could allow for various scanning (e.g. medical, material science, geological) allowing (emission) CT scanning for 3D maps of $N_{2}$ excited atoms/molecules between them as in Fig. 15. To observe negative radiation pressure for non-visible spectrum, there could be used antennas with excited resonator, or maybe X-ray lamp, or electrons travelling e.g. in synchrotron/vaccum tube lamp as having potential to emit such photons.

### v13

@@ I Introduction
Figure 4: In commercially available e.g. erbium-doped optical fiber amplifiers (EDFA, or simpler semiconductor SOA, etc.) [11] there is used backward ASE (amplified spontaneous emission): increased emission from pumped doped fiber, also toward laser e.g. in [12], exactly as in the sought CPT analog of ”laser causes excitation” effect, usually removed by forward optical isolators. Top: Analogously to STED-like test, we can verify this photon direction by measuring delay between the impulse and reduced population of excited atoms of external target, e.g. by monitoring its spontaneous emission. For this purpose, the shown standard forward optical isolator should be removed, or rotated to backward position. Bottom: simplified setting from [12] seems much more practical for tests: of delay (opposite if CPT analog), separation (of forward/backward action), transparency (forward blocked by $N_{1}$, so backward should be by usually much lower $N_{2}$) - allowing for potential future applications. As in Fig. 5, it seems crucial to use reverse bias for the laser, in practice (e.g. [1, 2, 12, 13]) probably included in varying/impulse signal.
@@
Figure 5: Usually system has tendency to get to the ground state, emitting photon, which e.g. excites some atom, according to absorption equation. However, getting reversed tendency should increase amplitude (probability) of shown symmetric reversed Feynman diagram: causing absorption of photon, emitted by some atom - according to symmetric stimulated emission equation. Beside previous based on synchrotron radiation, diagram shows two examples of obtaining such revered tendency: upper uses membrane GFP (green fluorescent protein) which can act as light-driven protein pump [18], in reversed proton gradient would get tendency to absorb photons, what might be used e.g. by biological organisms for thermoregulation. The bottom is analogously applied reversed bias/potential/electron gradient for light/laser diode, what is also applied for light detectors[19], but should actually also stimulate emission - e.g. used as absorber in [13] clearly helps with generation of photon impulse. Impulse laser sources often have e.g. sinc-like shape, so should also include reversed bias, what might explain observations in [1, 2]. Backward ASE in [12] probably also uses both forward and reverse bias in varying signal, if successful would be perfect for presented applications. Also e.g. Fig. 2, 4 tests should rather use reverse bias. Alternatively, change of $N_{2}$ above/below thermal $N_{2}\sim N_{1}\exp(-\beta(E_{2}-E_{1}))$ of Boltzmann distribution should also increase amplitude of Feynman diagram with emission/absorption there, and absortpion/emission in the target.
@@
Figure 7: STED microscopy (image source) is a popular application of stimulated emission - using additional laser reducing excitation in a doughnut shape, leaving smaller dot to increase resolution.
@@ II-B Alternative STED-like tests (needs reverse bias)
Figure 9: Ring laser uses optical isolator to enforce unidirectional photon circulation, which would be reversed in CPT perspective for reversed bias - switching the action of absorption and stimulated emission equations, suggesting only one of them should act on external targets - let us refer to them as forward/backward beam. Analogously for reverse biased lasers, or synchrotron sources like synchrotron, betatron, wiggler, undulator, free electron laser, bremsstrahlung - reversing electron and photon trajectory/tendency in CPT perspective. FEL reaches superradiance (power $P$ growing with square of the number of electrons $N_{e}^{2}$), better coherence, narrower spectrum and angle [16] - what already seems a consequence of discussed CPT analog of emission: that further electrons stimulate emission from earlier, as in T/CPT view there would be reversed photon exchange: further electrons exciting earlier.
Figure 10: General relativity below black hole horizon rotates time to space direction, in theory could also twice further e.g. in Klein-botle like nonorientable wormhole [28]: applying time symmetry to a rocket travelling through it, getting thought experiment valuable to gain intuitions about time symmetry. For example entropy gradient inside would be reversed. State preparation there would be postparation, allowing for 2WQC [29] more symmetric quantum computers (Fig. 14). Forward beam from such rocket would be backward beam for external observer, switching positive and negative radiation pressure. CT scanner there used on external object would map emission coefficient instead of absorption (Fig. 17). Our black holes for observer inside would be white holes (Fig. 19).
@@ IV Some potential new applications
The above considerations could lead to many new applications, especially if generating backward beam of negative radiation pressure acting on target only with stimulated emission equation. Optimistically it can be obtained by reverse biased diode/laser (without or with reversed optical isolator). Otherwise, as in Fig. 3, 9, should be possible with antennas and synchrotron radiation, e.g. betatron, synchrotron, free electron laser, or bremsstrahlung e.g. in X-ray lamp in strong magnetic field. For example:
@@
    Figure 18: Potential applications of proposed backward camera to obtain complementary information from astronomical objects, e.g. might allow detection/observation of black holes as in Fig. 19, 22, 23, maybe also finding some novel currently unexpected objects, and mapping emission asymmetry for cosmological studies. Many of such objects e.g. pulsars (image source) have accelerating charged particles, what leads to synchrotron radiation - e.g. causing excitation in cooled detector of standard telescopes. In CPT perspective of this scenario, there is still accelerating charge (in opposite direction) - should emit photons, causing excitation now toward our negative time, which is deexcitation toward positive time. Therefore, such process required by CPT symmetry would need excited telescope sensor - usually they are cooled instead, what might be the reason for not observing such effect, could lead to worth searching consequences like cooled halos around such pulsars. To have a chance to observe such effect, we would need to pump the sensor/detector (instead of standard cooling), again monitoring its population, or maybe using semi-transparent mirror as in Fig. 15 with various photon sources (e.g. antenna, synchrotron, X-ray lamp, forward bias diode with monitored current), maybe also include some optical isolator to allow only outgoing photons. If successful, such telescope should see (reversed) synchrotron, but not thermal radiation - e.g. allowing to separate them when combining with standard picture, also for observations/diagnostics/early warnings for our Sun [53].
@@
    Figure 19: Applying T/CPT symmetry to our black hole (running movie backward), it is a white hole, and they actively emit - acting with positive radiation pressure $\vec{p}=\langle\vec{E}\times\vec{H}\rangle/c$, heating matter around, what returning to our time perspective means that black hole should also actively cool matter around with negative radiation pressure. Hence white/black holes should repel/attract not only gravitationally, but also EM by radiation pressure - continuously interacting/exchanging information with the surrounding in symmetric way, this way allowing unitary evolution, avoiding information paradox [54] symmetrically for both white and black hole. It also would help forming dynamical outward/inward jets for both, and indeed they turn out very dynamical [55]. As white holes would heat matter around, black hole should symmetrically cool it - with the same spectrum, but replacing absorption with stimulated emission equation, and indeed matter [56], especially electrons [57] around turn out colder than expected, it also might prevent observation of IMBH (intermediate mass black holes). Such pull not only in gravity, but also in EM, would additionally increase growth of black hole, which indeed turns our faster than expected [58], generally might be faster than Hawking radiation - suggesting to reconsider their future evolution. While we might already observe indirect consequences of such stimulated emission caused by black hole, backward telescope like in Fig. 15, 18, 22 might also observe it directly - also below horizon, symmetrically to white hole observation. Black holes are hypothesized to have negative pressure [59], negative temperature pressure [60], which should also lead to radiation pressure through reversed emission/absorbtion tendency as in Fig. 5.
@@
    Figure 22: Potential observation of black holes with (radio)telescope. From CPT perspective: running movie backward, this would be white hole emitting photons (from below horizon) which could be absorbed by telescope - what in our time perspective means emission by telescope stimulated by black hole, however, it requires earlier excitation of telescope sensor. While usually radio telescope amplifiers are cooled, avoiding it or actively exciting resonator inside, in theory this antenna could also emit directionally. However, used frequencies are often difficult to absorb in cosmos, without specific absorber might go to infinity, what in CPT perspective would mean such EM wave was never emitted, hence emission should be forbidden if not pointing absorber, like black hole - suggesting novel way to detect them, and observe also below horizon, symmetrically to white hole emitting from below horizon. Another view is by $\langle\psi_{f}|U|\psi_{i}\rangle$ S-matrix: with amplitude/probability of photon depending both on emitter in $\psi_{i}$, but also absorber in $\psi_{f}$, they are switched in CPT view - without absorber, in CPT view there is no emitter. In contrast, LIGO measures lengths, which are T/CPT symmetric: should be able to see both retarded and advanced as in Fig. 23.
    Figure 23: Wheeler-Feynman absorber theory requires both standard retarded EM waves, but also symmetrically advanced, and they are switched in T/CPT view. As general relativity is rather solved by the least action principle, it should be also true for gravitational waves. LIGO measures lengths, which are T/CPT invariant, so should see both retarded and advanced waves - the big question is if there exist events for the latter. Top: observed waves from e.g. black hole mergers come mainly from orbiting of two masses, what is practically the same from time symmetry view - so should emit gravitational waves in both perspectives. Using forward Euler-Lagrange evolution would give retarded waves, but we could also use it backward, or should rather solve with the least action principle - both rather requiring also symmetric advanced wave, of similar chirp shapes at least in orbiting regime. As in the list of gravitational wave observations there was still observed only single EM counterpart from around 300 observed events, $\approx 10$ involving neutron stars, some of them could turn out advanced waves, especially that there should be more opportunities in our future than our past. Some suggestions for advanced waves could be events claimed to happen earlier than possible like GW190521 [69], or extremely large luminosity distances - now observed up to twice the age of the Universe. Additionally, MeerKAT pulsar timing array observe vibrations of the Universe [70] which seem to require more orbiting supermassive black holes than expected - including advanced could help. Bottom: another hypothetical scenario, as in Fig. 20 assuming Big Crunch - from which backward evolution seems similar to forward from Big Bang, hence both should have tendency to form black holes, which mergers should have reversed shape chirps - suggesting to search LIGO historical data for dataset of chirp templates with applied time symmetry. Another candidate for advanced wave might core-collapse events, forming pulling wave.
@@
  • Backward camera as in Fig. 15 maybe exploiting transparency and/or opposite delay e.g. for astronomical observations (of e.g. synchrotron radiation of pulsars, our Sun [53], maybe black holes as in Fig. 18, 19)- focused on target’s negative radiation pressure (or transmission of added backward beam), while standard (forward) camera is focused on positive. It would replace CMOS/CCD matrix behind the lens, with matrix/surface of continuously excited e.g. approximately 3-level medium like dye or erbium, plus some monitoring of its population map - e.g. through spontaneous emission, or EIT (electromagnetically induced transparency), half-silvered mirror and CMOS/CCD, or maybe forward biased diode with monitored current. Pairing with backward beam could allow for various scanning (e.g. medical, material science, geological) allowing (emission) CT scanning for 3D maps of $N_{2}$ excited atoms/molecules between them as in Fig. 17. To observe negative radiation pressure for non-visible spectrum, there could be used antennas with excited resonator, or maybe X-ray lamp, or electrons travelling e.g. in synchrotron/vaccum tube lamp as having potential to emit such photons.